OpenStudy (anonymous):

You are planning to use a ceramic tile design in your new bathroom. The tiles are blue-and-white equilateral triangles. You decide to arrange the blue tiles in a hexagonal shape as shown. If the side of each tile measures 4 centimeters, what will be the exact area of each hexagonal shape?

OpenStudy (anonymous):

The area of a equilateral triangle is \[A = \frac{ \sqrt{3} }{ 4 }a^{2}\] If we use 6 triangles the area will be \[6A = \frac{6 \sqrt{3} }{ 4 }a^{2}\] Now input 4cm for a

OpenStudy (anonymous):

i got 41.56922

OpenStudy (anonymous):

That is correct, well done!

OpenStudy (anonymous):

thank you so much!!